OpenStudy (anonymous):

Suppose Q = x^2 y^1 . If dQ/dt = -4 and dx/dt= -3, find dy/dt when x = 4 and y = 1 .

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get for dQ/dt in general?

OpenStudy (anonymous):

I believe this is an implicit differentiation problem

OpenStudy (anonymous):

Everything is a function of t. x and y are not special.

OpenStudy (anonymous):

for dq/dt i got 2x2yy'

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use the product rule to get: \[\Large Q = x^2y^1\] \[\Large \frac{dQ}{dt} = 2x*y*\frac{dx}{dt}+x^2*\frac{dy}{dt}\] Now plug in the given info and simplify.

OpenStudy (anonymous):

okay, got it, thanks

jimthompson5910 (jim_thompson5910):

np